We sampled 40 NY teachers for their annual income. In our sample,
the average annual income is $60,000 and the standard deviation of annual
incomes is $7,500. Based on that sample, we want to predict the mean annual
income μ of all the NY teachers with the confidence level of 0.99 (99 percent).
Part a) What is the error bound/margin E in our prediction about μ?
Part b) What prediction do we make about μ?

Answer by Theo(13342)

the sample size is 40
the sample mean is 60000.
the sample standard deviation is 7500.

you would use the t-score to help solve this.
the formula is t = (x-m)/s
t is the t-score
x is the sample mean
m is the population mean
s is the standard error.

the standard error is equal to the standard deviation / square root of sample size = 7500 / sqrt(40) = 1185.854123.

at 99% confidence interval, you would calculate the two-tailed critical t-score with 39 degrees of freedom (sample size minus 1) to be plus or minus 2.707913179.

on the high side of the confidence interval, your t-score formula becomes:
2.707913179 = (60000 - m) / 1185.854123.
your margin of error on the high side is equal to (60000 - m) which is equal to 2.707913179 * 1185.854123 = 3211.190008.
this indicates that your mean is 60000 - 3211.190008 = 56788.80999.

your solution should be:

Part a) What is the error bound/margin E in our prediction about μ?
margin of error = 3211.19 rounded to 2 decimal places.

Part b) What prediction do we make about μ?
population mean is assumed to be 56788.81 rounded to 2 decimal places.